An author has prepared to submit six books for publication. The probability of any book being accepted is 0.20, assuming independence, find the probability that the author will have:


(i) Exactly one book accepted.


(ii) At least two books accepted.


(iii) At most two books accepted.


1
Expert's answer
2021-01-26T04:07:07-0500

This is a binomial distribution with n=6, p=0.2.

(i) "P(X=1)=C_6^10.2^1(1-0.2)^5=0.3932."


(ii) "P(X\ge2)=1-P(X=0)-P(X=1)=C_6^00.2^0(1-0.2)^6+C_6^10.2^1(1-0.2)^5="

"=0.3446."


(iii) "P(X\le2)=P(X=0)+P(X=1)+P(X=2)="

"=C_6^00.2^0(1-0.2)^6+C_6^10.2^1(1-0.2)^5+C_6^20.2^2(1-0.2)^4=0.9011."